# Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities

### v1

@@ 7 Conclusion
In this paper, we introduced and explored Community-Cross-Instruct, a fully automated framework to represent online communities and elicit their views. By automating the process of surveying diverse online communities, Community-Cross-Instruct offers a cost-effective and efficient alternative to traditional survey methods used by social scientists. Through finetuning LLMs based on community-specific instructions and answers, Community-Cross-Instruct enables the generation of accurate responses that align with the beliefs and perspectives of different online communities. This innovative framework has demonstrated its effectiveness in representing various communities, including political and fitness forums on Reddit. Community-Cross-Instruct opens up new possibilities for researchers to gain insights into the diverse views and opinions present in online communities, paving the way for deeper understanding and analysis of digital societies. Moving forward, we will apply the framework to model organic online communities, such as those in the retweet network.
@@
We use Claude-3 for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as Claude-3. This reliance on a single LLM introduces a limitation, as it may lead to a bias where the evaluation dataset (CommSurvey) is inherently aligned with the model used for instruction generation (CommInst), potentially overestimating the alignment accuracy of the finetuned models. To address this limitation, we plan to incorporate a diverse set of advanced LLMs as they become available in future iterations of our framework. Additionally, we will include human-in-the-loop validation to ensure the generated datasets maintain high quality and representativeness, mitigating any biases introduced by the single LLM dependency. Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring robustness and generalizability of our results beyond the influence of a single LLM.
@@
To approximate group-level behavior, community members represented in the minority are inherently excluded. Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for bimodal distribution. We hope to build on this work by experimenting with survey designs that account for more diversity in communities.
@@
LLM responses are sensitive to slight changes in prompts Salinas and Morstatter (2024). In this work, we work primarily with one prompt for instruction generation. We would be interested to see how model’s generated instructions shift with different prompting schemas.
@@
Users might not be informed of how their data (reflected in their posts) are used to produce digital twins that mirror their voices and the purpose of the survey constructed based on their data. Furthermore, automatically surveying communities can reveal unconsented insights of certain individuals or groups of people online. To address these ethical concerns, we will implement several measures. Firstly, we will ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy. Secondly, we will seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure. Thirdly, we will obtain explicit consent from users where possible, clearly communicating the purpose and scope of our research. Lastly, we will adhere to ethical guidelines and institutional review board (IRB) requirements to ensure that our methods respect the privacy and rights of the individuals and communities involved.
Appendix A Data Details
@@ A.1 Subreddits
r/Anarchism
r/Anarchism is a forum with 268K members for discussing anarchism, assuming
participants are anarchists, allies, or learners.
@@
r/Anarcho_Capitalism is a subreddit with 197K members dedicated to discussing
@@
r/Conservative, with 1.1M members, offers the largest space on Reddit for fiscal
and social conservatives to explore and discuss political and cultural issues from
a distinctly conservative perspective.
r/Libertarian
r/Libertarian is a forum with 506K members focused on libertarianism and the
principles of a free society, explicitly rejecting anti-libertarian ideologies
like communism.
r/Republican
r/Republican is a partisan subreddit with 191K members specifically designed
as a space for Republicans to discuss issues among themselves.
r/democrats
r/democrats is a subreddit with 438K members dedicated to supporting the
Democratic Party, offering daily updates, policy analysis, and opportunities to
help elect Democrats at all levels.
r/Fitness
r/Fitness is the largest community dedicated to achieving physical fitness
goals with 12M members
r/loseit
r/loseit is a supportive community with 4M members for individuals of all
sizes to discuss healthy and sustainable weight loss methods
@@
r/keto is a community of 3.6M members for sharing experiences and advice
about the low-carb Ketogenic diet, which supports a range of health
conditions from diabetes to epilepsy
r/xxfitness
r/xxfitness is a community of 2.2M members for female and gender
non-binary/gender non-conforming Redditors to discuss fitness
r/GettingShredded
r/GettingShredded is a community of 162K members for anyone aiming to lose
body fat, improve their physique, and boost confidence through healthy fat loss
and strength training discussions.
Table 2: Reddit forums used in this study. Descriptions of (top) politics and (bottom) fitness subreddits.
A.2 Statistics of CommInst
Table 3 describes the number of demonstrations in each subreddit. The variance is due to their different topic coverage. Figure 6 presents the distribution of instruction lengths and answer lengths in CommInst-Pol and CommInst-Fit.
Politics
r/Anarch. r/Anarcho. r/Conserv. r/Democr. r/Libert. r/Republic.
Comments 182,192 589,354 2,391,745 142,274 1,950,089 140,572
Submissions 31,567 41,264 356,842 90,001 40,967 48,808
Demonstrations 2,430 3,027 2,943 3,159 2,928 2,574
Fitness
r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
Comments 907,839 967,645 603,920 420,082 78,053
Submissions 2,430 3,027 2,943 3,159 2,928
Demonstrations 3,438 2,793 1,716 2,883 1,215
Table 3: Dataset size measured by number of comments and submissions for each community in political and fitness communities, and the number of learned open-ended instruction-answer pairs (demonstrations).
(a) Instructions in CommInst-Pol
(b) Instructions in CommInst-Fit
(c) Answers in CommInst-Pol
(d) Instructions in CommInst-Fit
Figure 6: Distribution of instruction and answer lengths in CommInst-Pol and CommInst-Fit.
Appendix B Implementation Details
@@ B.1 Topic Modeling with BERTopic
Following the pipeline of BERTopic Grootendorst (2022), we first obtain the text embeddings using all-mpnet-base-v2 Song et al. (2020). The embedding dimensions are reduced with UMAP, with n_neighbors 15 and n_components 5. Then the embeddings are clustered using HDBSCAN, with min_cluster_size 50. After fitting the model, we remove texts that are not assigned any topics. Table 4 and Table 5 present the top-10 most frequent topics in politics and fitness.
idx Topic Keywords Frequency
1 covid, vaccinated, masks, vaccines, healthcare, deaths, flu, pandemic, coronavirus, cdc 48,787
2 abortion, fetus, abortions, roe, body, choice, unborn, rights, parenthood, cells 11,256
3 climate, change, solar, emissions, pipeline, fuels, coal, power, supply, grid 9,035
4 guns, firearms, shootings, firearm, amendment, laws, armed, rifles, militia, nra 8,833
5 police, officer, officers, gun, force, law, crime, video, training, situation 7,934
6 kamala, woman, pelosi, vp, hillary, president, herself, biden, party, with 7,879
7 racist, racism, whites, racists, racial, minorities, democrats, privilege, whiteness, discrimination 7,785
8 libertarian, libertarians, libertarianism, party, government, left, ideology, vote, political, conservative 6,666
9 anarchism, anarchy, anarchist, anarchists, anarcho, hierarchies, communism, ideology, government, system 6,311
10 biden, bernie, obama, candidate, voters, debate, election, democrats, hillary, sanders 5,768
@@
idx Topic Keywords Frequency
1 lbs, bmi, down, cw, gained, progress, healthy, tips, weeks, started 18,296
2 keto, ketosis, carbs, diet, 20g, protein, macros, started, calories, ketogenic 13,078
3 leggings, clothes, wear, pants, shorts, bras, shirts, clothing, dress, lululemon 7,063
4 motivation, journey, progress, yourself, change, motivated, success, start, proud, back 5,142
5 water, scale, weighing, retention, trend, normal, plateau, fluctuates, gain, weekly 4,880
6 sleep, morning, hours, work, gym, motivation, schedule, waking, exercise, routine 4,486
7 she, people, overweight, weight, obese, comment, healthy, skinny, talk, shaming 4,192
8 bf, percentage, fat, bodyfat, body, lean, bulk, estimates, lbs, pic 3,930
9 reps, sets, program, max, amrap, lifts, ranges, bench, 5x10, hypertrophy 3,788
10 shoulder, elbow, injury, arm, knees, physio, exercises, bicep, curls, heal 3,775
Table 5: Top-10 most frequent topics in fitness. Each topic is represented by the topic-10 keywords.
@@ B.2 Prompting Template for CommInst
The prompting template for generating the open-ended instruction-answer pairs in CommInst is shown in Figure 7.
@@
Figure 7: Template prompt to generate CommInst.
B.3 Prompting Template for Generating CommSurvey
The prompting template for generating the multiple choice instruction-answer pairs in CommSurvey is shown in Figure 8.
Figure 8: Template prompt to generate CommSurvey.
Appendix C Survey Sample
Samples of the surveys provided to human annotators for generating “semi-ground truth data” are provided in Figures 9 and 10 (political and fitness domains, respectively).
@@ Appendix C Survey Sample
Figure 9: Sample of the multiple choice survey generated for human evaluation in the politics domain.
@@
Figure 10: Sample of the multiple choice survey generated for human evaluation in the fitness domain.

### v2

@@ 8 Conclusion
In this paper, we introduced and explored Community-Cross-Instruct, a fully automated framework to represent online communities and elicit their views. By automating the process of surveying diverse online communities, Community-Cross-Instruct offers a cost-effective and efficient alternative to traditional survey methods used by social scientists. Through finetuning LLMs based on community-specific instructions and answers, Community-Cross-Instruct enables the generation of accurate responses that align with the beliefs and perspectives of different online communities. This innovative framework has demonstrated its effectiveness in representing various communities, including political and diet forums on Reddit. Community-Cross-Instruct opens up new possibilities for researchers to gain insights into online communities’ diverse views and opinions, paving the way for deeper understanding and analysis of digital societies.
Moving forward, we will apply the framework to model organic online communities, such as those in the retweet network. In addition, we are interested in aligning LLMs to communities through reinforcement learning from human feedback (RLHF).
@@
Ignoring Thread Structure in Reddit Data.
We treat each Reddit comment and submission as independent data points, without taking into account the thread structure in which they are embedded. Reddit discussions are inherently hierarchical, with comments often building upon or responding to previous comments and submissions. By ignoring this structure, we may lose important contextual information that could help capture the nuances of community interactions and opinions. Future work could explore leveraging the thread structure to better model the flow of conversations, capturing how community members engage with each other’s ideas and how opinions evolve within discussions.
@@
We use GPT-4o for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as GPT-4o. This reliance on a single LLM introduces a limitation, as it may lead to a bias where the evaluation dataset (CommSurvey) is inherently aligned with the model used for instruction generation (CommInst), potentially overestimating the alignment accuracy of the finetuned models. To address this limitation, we plan to incorporate a diverse set of advanced LLMs as they become available in future iterations of our framework. Additionally, we will include human-in-the-loop validation to ensure the generated datasets maintain high quality and representativeness, mitigating any biases introduced by the single LLM dependency. Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring the robustness and generalizability of our results beyond the influence of a single LLM.
@@
To approximate group-level behavior, community members represented in the minority might be inherently excluded. Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for a more complex distribution of the various opinions within the community. We hope to build on this work by experimenting with survey designs that account for more diversity in communities.
@@
LLM responses are sensitive to slight changes in prompts Salinas and Morstatter (2024). In this work, we work primarily with one prompt for instruction generation. We would be interested to see how the model’s generated instructions shift with different prompting schema.
@@
Finetuning LLMs towards Bias.
Aligning LLMs with specific communities may result in models that appear more biased, as they are finetuned to the distinctive views and perspectives of those communities. However, this process is done solely for the purpose of accurately reflecting the values and attitudes prevalent within each community, rather than intentionally making the models more biased. Our goal is not to reinforce or amplify harmful biases, but rather to provide a computational tool that can represent the views of a given community for research purposes. To mitigate potential misuse, these models should be used in controlled environments and in contexts where understanding specific community perspectives is essential for social, cultural, or political research.
User Privacy and Consent.
Users might not be informed of how their data (reflected in their posts) are used to produce digital twins that mirror their voices and the purpose of the survey constructed based on their data. Furthermore, automatically surveying communities can reveal unconsented insights of certain individuals or groups of people online. To address these ethical concerns, we implement several measures. Firstly, we ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy. Secondly, we seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure. Thirdly, we will obtain explicit consent from users where possible, clearly communicating the purpose and scope of our research. Lastly, we will adhere to ethical guidelines and institutional review board (IRB) requirements to ensure that our methods respect the privacy and rights of the individuals and communities involved.
Acknowledgements
We are grateful to Patrick Gerard, Ashwin Rao, and DJ Berry for helping review responses to surveys. This project was funded in part by the Defense Advanced Research Projects Agency (DARPA) under Agreement Nos. HR00112290021 and HR001121C0168.
Appendix A Reddit Data
@@ A.1 Subreddits
r/Liberal
r/Liberal is a subreddit dedicated to discussions and news from a liberal political
perspective, focusing on progressive policies, social justice, and left-leaning viewpoints.
r/NeutralPolitics
r/NeutralPolitics is a community dedicated to evenhanded, empirical discussion of
political issues. It is a space to discuss policy and the tone of political debate.
@@
r/Anarcho_Capitalism is a subreddit dedicated to discussing
@@
r/Conservative offers the largest space on Reddit for fiscal and social
conservatives to explore and discuss political and cultural issues from a distinctly
conservative perspective.
r/AskThe_Donald
r/AskThe_Donald is a subreddit that serves as a hub for supporters of former President
Donald Trump, fostering discussions around conservative politics, pro-Trump views,
and right-wing ideologies.
@@
r/keto is a community for sharing experiences and advice about the low-carb Ketogenic
diet, which supports a range of health conditions from diabetes to epilepsy
r/WeightLossAdvice
r/WeightLossAdvice is a subreddit where users share tips, strategies, and support for healthy
and sustainable weight loss, with a focus on practical advice and personal experiences.
r/EDAnonymous
r/EDAnonymous is a subreddit that provides a supportive, anonymous space for individuals
struggling with eating disorders to share their experiences, seek advice, and offer
encouragement on the path to recovery.
Table 3: Reddit forums used in this study.
r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
# of comments 31,233 35,725 670,686 2,243,842 142,543
# of submissions 5,243 2,072 22,551 153,813 15,645
r/keto r/WLAdvice r/EDAnonymous
# of comments 603,466 225,635 347,477
# of submissions 49,571 42,840 99,925
Table 4: Number of comments and submissions in each subreddit.
Appendix B Instructional Data Generation
@@ B.1 Topic Modeling with BERTopic
Following the pipeline of BERTopic Grootendorst (2022), we first obtain the text embeddings using all-mpnet-base-v2 Song et al. (2020). The embedding dimensions are reduced with UMAP, with n_neighbors 15 and n_components 5. Then the embeddings are clustered using HDBSCAN, with min_cluster_size 40. After fitting the model, we remove texts that are not assigned any topics. Table 5 and Table 6 present the top 10 most frequent topics in politics and diet.
idx Topic Keywords
1 vaccine, covid, vaccinated, vaccines, flu, unvaccinated, vaccination, mrna, pandemic, data
2 climate, prices, solar, change, co2, emissions, coal, cars, fuels, earth
3 ballots, fraud, election, voter, ballot, evidence, voters, machines, audits, rigged
4 abortion, abortions, fetus, roe, murder, unborn, conception, choice, womb, cells
5 ukraine, russia, putin, nato, war, ukrainians, crimea, invade, conflict, eu
6 twitter, facebook, musk, google, social, platforms, companies, users, censorship, ban
7 gun, guns, firearms, shootings, firearm, rifles, laws, amendment, ammo, armed
8 impeachment, mueller, fbi, impeach, documents, collusion, comey, congress, whistleblower, crimes
9 gender, trans, transgender, lgbt, dysphoria, children, sexuality, identity, pronouns, genders
10 capitol, blm, antifa, riots, riot, insurrection, rioters, protesters, terrorism, january
@@
idx Topic Keywords
1 keto, protein, ketosis, started, macros, back, weeks, carbs, diet, calories
2 skinny, thin, girls, like, overweight, body, underweight, weight, hate, myself
3 thank, therapy, relapse, recover, help, life, support, yourself, proud, treatment
4 scale, week, water, lose, weighing, daily, trend, fluctuates, months, plateau
5 potassium, sodium, magnesium, electrolytes, mg, supplement, ketoade, citrate, 5000mg, chloride
6 alcohol, drinking, beer, drink, drinks, liquor, alcoholic, gin, booze, drinker
7 protein, macros, fat, min, kcal, 20g, carbs, bf, lean, need
8 binging, restricting, binges, eating, hunger, control, bingeing, cycle, guilt, feeling
9 ed, eds, recovery, recover, therapist, treatment, me, help, coping, talk
10 coffee, tea, chocolate, brew, creamer, starbucks, latte, unsweetened, flavors, sweetener
Table 6: Top-10 most frequent topics in diet. Each topic is represented by the topic-10 keywords.
@@ B.2 Prompting Template for CommInst
The prompting template for generating the open-ended instruction-response pairs in CommInst and multi-choice question-answer pairs in CommSurvey is shown in Figure 4.
@@
Figure 4: Prompting template to generate CommInst and CommSurvey in the politics domain.
B.3 Statistics of CommInst and CommSurvey
Table 7 describes the number of open-ended instruction-response pairs in CommInst and multi-choice question-answer pairs in CommSurvey. The variance is due to their different topic coverage.
r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
# of instructions 234 234 300 303 303
# of survey questions 156 156 200 202 202
r/keto r/WLAdvice r/EDAnonymous
# of instructions 1,032 885 870
# of survey questions 688 590 580
Table 7: Number of generated open-ended instruction-response pairs in CommInsti and multi-choice question-answer pairs in CommSurveyi, for each community.
B.4 Community Agreement Analysis
For a pair of communities $C_{i}$ and $C_{j}$, we identify their common survey questions in CommSurveyi and CommSurveyj, and compute their agreement on the questions using Cohen’s Kappa. The agreement matrices for politics and diet are shown in Figure 5.
In the politics domain, we observe significant polarization between communities. For example, the agreement between r/Liberal and both r/Conservative (0.01) and r/AskThe_Donald (0.03) is extremely low, indicating the deep ideological divide between these left-leaning and right-leaning communities. This low agreement reflects how vastly different their responses are to the same political survey questions, aligning with the broader patterns of political polarization seen across online platforms. On the other hand, r/Conservative and r/AskThe_Donald exhibit much higher agreement (0.59), highlighting the ideological overlap between these conservative-leaning communities, likely reflecting their shared values and viewpoints on key political issues.
Meanwhile, r/NeutralPolitics shows moderate agreement with both r/Anarcho_Capitalism (0.24) and r/Liberal (0.28), suggesting that as a centrist or neutral forum, it shares certain perspectives with both libertarian and liberal ideologies. Additionally, r/Anarcho_Capitalism shows higher agreement with conservative-leaning subreddits, such as r/Conservative (0.36) and r/AskThe_Donald (0.44), reflecting shared economic or libertarian principles, despite divergences on other political issues.
In the diet domain, the relationships between the communities are more practical and less polarized compared to politics. Communities like r/keto and r/WeightLossAdvice exhibit moderate agreement (0.39), likely reflecting shared goals related to weight loss and health, even though they may emphasize different strategies. Similarly, r/EDAnonymous and r/keto (0.38) show moderate overlap, suggesting that while their focus areas are different—one being centered on eating disorders and the other on ketogenic diets—they share some common ground in their approaches to health and diet-related topics.
The strongest alignment within the diet domain is between r/WeightLossAdvice and r/EDAnonymous (0.41), which likely stems from overlapping concerns about diet and body image, which are central to both communities’ discussions.
(a) politics
(b) diet
Figure 5: Pairwise agreement between different communities, measured by Cohen’s Kappa.
Appendix C Human Annotation
We verify the faithfulness of CommSurvey as the “semi-ground truths” for the communities via human annotation. We also tried directly posting the survey questions to the subreddits and having community members answer them. However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits. This further indicates the difficulty of directly surveying online communities.
C.1 Annotator Recruiting
The four annotators for the politics domain closely follow political news and are knowledgeable about the platforms of the parties represented in our sample. The three annotators for the diet domain are active members of diet and diet and have substantial knowledge of diet-related topics and trends. To ensure the annotators were capable of accurately answering the survey questions for different communities, we implemented a thorough selection and training process. Annotators were chosen based on their demonstrated expertise and familiarity with the relevant domains. Additionally, we conducted preliminary tests where annotators answered a subset of questions, and their responses were evaluated for consistency and accuracy.
C.2 Annotation Process
We randomly sampled 20 questions from CommSurvey-Politics (Figure 6) and 10 questions from CommSurvey-Diet (Figure 7) for the annotators to answer. When answering each question, the annotators were instructed to search the relevant posts from the subreddit and learn their views from the discussions, instead of solely relying on their pre-assumption about the community. Each annotator filled out the question for a different subset of the communities, as shown in Table 8. If multiple annotators annotated for a community, they had another round of discussions to resolve their disagreement. As a result, for community, the annotators delivered one set of annotations of the survey questions.
@@ C.2 Annotation Process
Figure 6: Ten multiple choice survey questions (out of the twenty) in CommSurvey for human evaluation in the politics domain.
@@
Figure 7: Ten multiple choice survey questions CommSurvey for human evaluation in the diet domain.
C.3 Annotation Evaluation
r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
Annotator 1 x x
Annotator 2 x x x x
Annotator 3 x x
Annotator 4 x x
r/keto r/WLAdvice r/EDAnonymous
Annotator 1 x x x
Annotator 2 x x
Annotator 3 x x
Table 8: Communities that each annotator annotated for.
For each subreddit, we compare the LLM-generated “semi-ground truths” to the label from human annotators, and present the accuracy in Table 9. In both domains, “semi-ground truths” achieve strong agreement with human annotations for most subreddits, which gives confidence that the advanced LLM-generated survey answers can be used as “semi-ground truth” to evaluate finetuned foundational LLMs.
r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
Human-LLM Acc 0.75 NA 0.65 0.55 0.55
r/keto r/WLAdvice r/EDAnonymous
Human-LLM Acc 0.7 0.6 0.5
Table 9: Agreement between human annotators’ survey responses and the advanced LLM (GPT-4o) generated survey answers, measured by accuracy.